### v1

@@ Abstract
We observed comet C/2025 A6 with FAST telescope equipped with the ultra-wideband receiver from 23rd October to 8th November 2025 and detected the OH 18-cm lines for the first time. The OH lines underwent a reversal from emission to absorption from 23rd October to 5th November, which is mainly caused by variations in the heliocentric velocity. Through trapezoidal fitting of the OH line profiles, we derive expansion velocities of the water that rise as the heliocentric distance decreases. Based on these results, we estimated the OH production rates of C/2025 A6 for 23rd October, 26th October, 4th November, and 5th November and it presents a significant upward trend.
@@
The comets spend most of their lives in the outer, frozen regions of the solar system. Their chemical compositions preserve crucial information about the primordial environment and the processes that may dominated the Solar System’s formation (2017RSPTA.37560261A; 1991ASSL..167.....N). When comets approach the Sun, the solar radiation heats their nuclei, causing subsurface ices to sublimate, and releasing gases and various molecules into the space. As the primary volatile in comets, the production rate and abundance of water provide directly insights into the activity of the nucleus and its physical properties (2004come.book..391B). Under the solar radiation, the water molecules undergo photodissociation: + H2O , producing hydroxyl radicals (OH, 1989A&A...213..459C).
@@
Hydroxyl radicals in cometary coma are primarily observed at ultraviolet and radio wavelengths. At radio frequencies, the OH 18-cm lines arise from -doubling and hyperfine transitions within the ground state of the OH radical, distributed near 1612, 1665, 1667, and 1720 MHz (see Table 1). Compared with UV measurements, OH 18-cm lines can be observed from the ground, and less affected by dust extinction. Their narrow profiles provide direct constraints on the gas expansion velocity and kinematic structure of the coma (1990A&A...238..382B). The OH 18-cm line profiles can be used to measure the expansion velocity of water in the cometary coma (1990A&A...238..382B), providing a basis for further studies of the coma’s gas dynamics. Since the first detection the OH 18-cm lines from Comet Kohoutek in 1973 (1974A&A....34..163B), long-term ground-based observations have established extensive datasets of water and OH emissions for a wide variety of comets (e.g., 1981A&A....99..320D; 2002A&A...393.1053C; 2017AJ....154..249W). These studies have extended observations to comets with different orbital periods and have used variations in molecular production rates to investigate cometary outgassing activity. On the other hand, either the emission or the absorption of the OH 18-cm lines from a comet could be observed, where OH is produced by the reaction between the photon and water molecular, requiring the the ultraviolet light with the wavelength shorter than 186 nm (1989A&A...213..459C).
The Comet C/2025 A6 (Lemmon) is a non-periodic comet. It was discovered by the Mount Lemmon Survey on 3rd January, 2025, and passed closest to Earth on 21st October, 2025, with a geocentric distance of 0.59 AU. It subsequently reached perihelion on 8th November, 2025, at a heliocentric distance of 0.53 AU. Between 3rd January, 2025 and 11th November, 2025, it is in the sky coverage of the Five-hundred-meter Aperture Spherical radio telescope (FAST), which is the largest single-dish radio telescope nowadays in the world. The FAST has been used for comet related studies before (e.g., 2024RAA....24j5008C) with the L-band to search for the molecular spectral lines.
@@ 2 Observations
The observations of C/2025 A6 were performed with FAST (2011IJMPD..20..989N) on 23rd and 26th October and from 3rd to 8th November in this year. The details can be found in Table 2. It is worth noting that the C/2025 A6 moved very quickly, all observations were conducted in the user-defined mode.
In order to observe the OH 18-cm lines at 1700 MHz, the ultra-wideband (UWB) receiver (2023RAA....23g5016Z) covering 0.5–3.3 GHz was used. This testing receiver is not cooled and with a system temperature between 89 to 130 K. At the spectra frequency (1660 MHz), the beam size is around . A photo of the comet with the size of FAST beam showing that the nuclei can be fully covered (Figure 1). Previous tests have shown excellent agreement between the OH flux densities and velocities measured with the FAST UWB receiver and those obtained with the Arecibo 300 m telescope (2023RAA....23g5016Z).
@@
| Start Time (UTC) | R.A. (h:m:s) | Dec. (d:m:s) | (AU) | (km s-1) | (AU) | (km s-1) | (K Jy-1) | Total times (s) | Trajectory |
@@ 3 Data reduction
Due to the limitation of sideband suppression of the UWB, only the bands 500–950, 950–1750, 1750–2550, and 2550–3300 MHz were used (2023RAA....23g5016Z). The OH 18-cm lines are around 1612 to 1720 MHz. Thus, only the spectra in such frequency range were used in our study. For each cycle, the temperature corresponding to the measured power is determined by differencing the spectral line data acquired with open and closed noise diode, followed by calibrating the resulting power values of all spectral lines into temperature. The calibrated spectral data are processed using the astronomical data processing software GILDAS/CLASS111https://www.iram.fr/IRAMFR/GILDAS/. For all observations, a second-order polynomial baseline was fitted and subtracted over a velocity span of 60 km s-1 around the line, excluding the channels containing line emission, absorption features, or identifiable RFI.
@@
We performed Gaussian fitting on all spectra with clear detections to obtain the integrated intensity, line-center velocity, FWHM, and peak amplitude of the OH 18-cm lines. We also applied a trapezoidal fitting procedure to the 1667 MHz spectra of the four observations with detections (23rd October, 26th October, 4th November and 5th November) in order to derive the half-width of the trapezoidal base. The derived parameters are summarized in Table 3.
@@
| Date (UT) | Lines | RMS (K) | Area (K km s-1) | (K) | S ( Jy km s-1) | (km s-1) | FWHM (km s-1) | (K) | (km s-1) | |
| 2025/10/23 | 1665 | 0.07 | 2.10.1 | 0.45 | 3.3 | 1.70.1 | 13.40.1 | 3.40.2 | 0.59 | 2.40.3 |
| 1667 | 0.07 | 3.90.1 | 3.10.1 | 13.30.1 | 3.10.1 | 1.21 | 2.40.1 | |||
| 2025/10/26 | 1665 | 0.06 | 2.30.1 | 0.51 | 3.5 | 1.80.1 | 28.80.1 | 3.50.2 | 0.61 | 2.90.6 |
| 1667 | 0.06 | 4.80.1 | 3.80.1 | 29.00.1 | 3.40.1 | 1.33 | 2.60.1 | |||
| 2025/11/03 | 1665 | 0.08 | -0.20.1 | -0.27 | 3.1 | -0.20.1 | 54.00.1 | 0.70.3 | -0.23 | - |
| 1667 | 0.09 | -0.40.1 | -0.40.1 | 53.50.1 | 1.10.2 | -0.34 | - | |||
| 2025/11/04 | 1665 | 0.12 | -0.50.1 | -0.37 | 3.1 | -0.50.1 | 55.90.2 | 1.30.3 | -0.38 | - |
| 1667 | 0.09 | -1.40.2 | -1.30.2 | 56.40.2 | 3.20.4 | -0.40 | 3.60.8 | |||
| 2025/11/05 | 1665 | 0.20 | -0.60.2 | -0.33 | 3.1 | -0.70.2 | 57.80.6 | 3.91.1 | -0.16 | - |
| 1667 | 0.09 | -1.30.2 | -1.40.2 | 57.00.3 | 3.90.6 | -0.33 | 3.90.9 |
Notes. Column 2 represents the spectral lines observed in the corresponding frequency range; Column 3 is the 1 noise of the observed spectra with the channel width 0.38 km s-1; Column 4, 8, 9, 10 are derived from Gaussian fitting, representing the integrated intensity, the position of the center of the line obtained through fitting, the full width at half height, and the peak temperature obtained through fitting, respectively; Column 5 is the inversion of ground state of (1988ApJ...331.1058S); Column 7 represents the corresponding flux; Column 11 is the half-length of the base, derived from trapezoidal fitting which represents the sum of the velocities of the parent and daughter molecules.
@@ 4 Results and Discussion
During four observing epochs on 23rd October, 26th October, 4th November, and 5th November, either emission or absorption of the 1667 MHz line was detected.
@@ 4.1 Change of OH 18-cm Lines from Emission to Absorption
In our observations, the OH 18-cm lines exhibit a clear evolution from emission to absorption. During the first two observing epochs, the OH lines are detected in emission in both the 1665 and 1667 MHz transitions, with corresponding velocities of 13.4 km s-1 and 28.9 km s-1, respectively. In subsequent observations, the lines appear in absorption in both main lines, with a stronger signal at 1667 MHz. The corresponding line-of-sight velocities are 54.0 km s-1, 56.3 km s-1, and 57.0 km s-1 on 3th November, 4th November and 5th November, respectively. The fitted line-center velocities are consistent with the expected radial velocity of the comet relative to the telescope at each observing epoch, confirming the cometary origin of the detected signals. No distinct OH signal is detected during the observations from 6th to 8th November within the expected velocity range.
@@
such that at Å, a heliocentric radial velocity of corresponds to a wavelength of Å. At these wavelengths, the solar Fraunhofer spectrum is strongly structured on scales of order 0.1 Å, comparable to the Doppler shifts induced by heliocentric radial velocities of a few km s-1 (see Fig. 2 of 1982ApJ...258..864S).
Consequently, modest heliocentric radial velocity variations during the observing period can alter the OH pumping rate and drive the inversion parameter of the ground-state -doublet across zero. Against the Galactic background continuum, this leads to a transition in the observed OH 18-cm line profiles from emission to absorption (1981A&A....99..320D). Using the inversion curves of 1988ApJ...331.1058S, we derived the inversion parameter at the heliocentric radial velocity of each epoch and compared it with the observed line profiles. The emission detected at early epochs and the absorption observed at later times are consistent with positive () and negative () inversion, respectively. The absence of a clear OH detection in the final observing epochs likely indicates that the inversion parameter was close to zero.
@@ 4.2 Water Expansion Velocities from OH Lines
The OH coma is commonly described using a spherically symmetric, steady-state, and isotropic outflow model, in which OH is treated as a secondary product of H2O photodissociation. The velocity of OH molecules is determined by the expansion velocity of the parent gas, , combined with an additional ejection velocity due to photodissociation, (1984A&A...131..111B; 2002A&A...393.1053C). The OH 18-cm lines profiles can be well reproduced by trapezoidal fitting, where the width of the lower base corresponds to 2(+). This trapezoidal profile model has been widely applied in cometary OH line studies (2023A&A...677A.157D; 2025A&A...701A.204L; 2025PSJ.....6..261S). Although OH line profiles always display asymmetries due to anisotropic outgassing toward the Sun or Earth (2025A&A...701A.204L), such effects have only a limited impact on the determination of the lower-base width (1984A&A...131..111B). In this work, we adopt a value of = 0.95 km s-1 to represent the ejection velocity of OH radicals following the photodissociation of H2O, based on the analyses presented in (2002A&A...393.1053C; 2007A&A...467..729T). This value is consistent with the typical OH ejection speeds reported in previous cometary studies (1989A&A...213..459C; 1990A&A...238..382B). The resulting values of derived from the symmetric trapezoidal fits are listed in Table 3.
@@ 4.3 The estimate of OH production rate
where is total number of OH radicals in the coma, is the integrated intensity of the 1667 MHz line in units of Jy km s-1, is the inversion of the ground state, and is the geocentric distance of the comet in AU. denotes the background temperature in the unit of K, is the OH lifetime, and is correction factor. For , it can be expressed as: = (2017AJ....154..249W; 2025PSJ.....6..261S), and was estimated as s from 1990A&A...238..382B, where the is the heliocentric distance of the comet. Following the method of 2002A&A...393.1053C, we estimated using the 1420 MHz brightness temperature distribution. The 1420 MHz map was obtained from the NASA/LAMBDA222https://lambda.gsfc.nasa.gov/product Stockert–Villa-Elisa all-sky survey (1982A&AS...48..219R; 1986A&AS...63..205R; 2001A&A...368.1123T), and the corresponding background temperature is listed in Table 3. The inversion of ground state at the corresponding coma velocity is obtained from (1988ApJ...331.1058S) and listed in Table 3.
@@
where is the OH column density distribution derived from Haser model. The expansion velocity is derived from the trapezoidal fits, and the photodissociation lifetime of H2O at 1 AU is adopted as s (1992Ap&SS.195....1H), also scaling as . The beam weighting function is assumed to be Gaussian, with a FWHM of 2.5 arcmin.
@@
When only the beam correction is applied, the derived values serve as lower limits, with value of , , , and s-1 for 23rd October, 26th October, 4th November, and 5th November, respectively. The production rate basically increases as the heliocentric distance decreases. The water production rate of comet can be approximated as (e.g., 2025PSJ.....6..261S; 2025A&A...701A.204L).

### v2

@@ Abstract
We observed comet C/2025 A6 with the Five-hundred-meter Aperture Spherical radio Telescope (FAST) equipped with the ultra-wideband receiver from 2025 October 23 to November 8, and it was the first detection for this comet with FAST. Through trapezoidal fitting of the OH line profiles, we derived the expansion velocities of the water which showed an increase from 1.50.3 km s-1 at the heliocentric distance of 0.65 AU to 3.00.9 km s-1 at 0.54 AU. Based on these results, we estimated the OH production rates of C/2025 A6 for October 23, October 26, November 4 and November 5 which were , , , and s-1 respectively. The results show a significant upward trend.
@@
The comets spend most of their lives in the outer, frozen regions of the solar system. Their chemical compositions preserve crucial information about the primordial environment and the processes that may dominated the Solar System’s formation (Newburn et al. 1991; A’Hearn 2017). When comets approach the Sun, the solar radiation heats their nuclei, causing ices to sublimate, and releasing gases and various molecules into space. Cometary activity can be probed across multiple wavelength regimes with ground-based telescopes, including optical and millimeter/radio observations, among others (e.g., Betzler 2024; Manna et al. 2024; Zhang et al. 2024). As the primary volatile in comets, the production rate and abundance of water provide directly insights into the activity of the nucleus and its physical properties (Bockelée-Morvan et al. 2004). Under the solar radiation, water molecules undergo photodissociation: + H2O , producing hydroxyl radicals (OH) via ultraviolet radiation at wavelengths shorter than 186 nm (Crovisier 1989).
@@
Hydroxyl radicals in cometary coma are primarily observed at ultraviolet and radio wavelengths. At radio frequencies, the OH 18-cm lines arise from -doubling and hyperfine transitions within the ground state of the OH radical, distributed near 1612, 1665, 1667, and 1720 MHz (see Table 1). Compared with UV measurements, OH 18-cm lines can be observed from the ground, and less affected by dust extinction. Their narrow profiles provide direct constraints on the gas expansion velocity and kinematic structure of the coma (Bockelée-Morvan et al. 1990). The OH 18-cm line profiles can be used to measure the expansion velocity of water in the cometary coma (Bockelée-Morvan et al. 1990), providing a basis for further studies of the coma’s gas dynamics. Since the first detection the OH 18-cm lines from Comet Kohoutek in 1973 (Biraud et al. 1974), long-term ground-based observations have established extensive datasets of water and OH emissions for a wide variety of comets (e.g., Despois et al. 1981; Crovisier et al. 2002; Wang et al. 2017). These studies have extended observations to comets with different orbital periods and have used variations in molecular production rates to investigate cometary outgassing activity. For comets, either emission or absorption of the OH 18-cm lines could be observed, mainly depending on the comet’s heliocentric radial velocity.
The Comet C/2025 A6 (Lemmon) is a long-period comet. It was discovered by the Mount Lemmon Survey on 2025 January 3 (Fuls et al. 2025), and passed closest to Earth on 2025 October 21, with a geocentric distance of 0.59 AU (from JPL horizons system111https://ssd.jpl.nasa.gov/horizons/, hereinafter). It subsequently reached perihelion on 2025 November 8, at a heliocentric distance of 0.53 AU. Since the semi-major axis of C/2025 A6 is only about 120 AU, which is considerably smaller than the 10,000 AU threshold typically used to define dynamically new comets (Levison 1996), the comet is likely a returning long-period comet rather than a dynamically new one. Comet C/2025 A6 provides an opportunity to measure water-driven activity in a long-period comet. The OH 18-cm emission from comets is intrinsically weak and strongly dependent on solar pumping conditions, so detecting it requires high-sensitivity single-dish observations. Between January 3 and November 11, 2025, it was observable with FAST, the largest single-dish radio telescope in the world. FAST’s enormous collecting area is particularly well suited for this purpose and enables the detection of OH emission from comets. Observations of C/2025 A6 therefore help expand the currently sparse dataset of OH 18-cm measurements and provide an important benchmark for studies of cometary water production rates. In addition, a larger aperture results in a smaller beam size, providing an opportunity to measure OH line intensities at different positions within the coma. This capability provides a useful complement to observations with smaller telescopes, whose beams often encompass the entire coma and therefore measure only the beam-averaged emission. The FAST has been used for comet related studies before (e.g., Chen et al. 2024) with the L-band to search for the molecular spectral lines.
@@ 2 Observations
The observations of C/2025 A6 were performed with FAST (Nan et al. 2011) on October 23, 26 and from November 3 to 8 in 2025. The details can be found in Table 2. It is worth noting that the C/2025 A6 moved very quickly, all observations were conducted in the user-defined mode.
In order to observe the OH 18-cm lines at 1700 MHz, the ultra-wideband (UWB) receiver (Zhang et al. 2023) covering 0.5–3.3 GHz was used. This testing receiver is not cooled and with a system temperature between 89 to 130 K. At the spectral frequency (1660 MHz), the beam size is around . An optical image of the comet, taken with the 70-cm UCASST telescope, showing that a significant fraction of the coma is covered (Figure 1). Previous tests have shown excellent agreement between the OH flux densities and velocities measured with the FAST UWB receiver and those obtained with the Arecibo 300 m telescope (Zhang et al. 2023).
@@
| Start Time (UTC) | R.A. (h:m:s) | Dec. (d:m:s) | (AU) | (km s-1) | (AU) | (km s-1) | (K Jy-1) | Total time (s) | Tracking mode |
@@ 3 Data reduction
Due to the limitation of sideband suppression of the UWB, only the bands 500–950, 950–1750, 1750–2550, and 2550–3300 MHz were used (Zhang et al. 2023). We searched for the CH lines in the 500–950 MHz sub-band. However, as the UWB receiver is still under commissioning, this band is significantly affected by radio frequency interference (RFI), and no CH line was detected. The OH 18-cm lines are located within the 800-1900 MHz sub-band, which is comparatively less affected by RFI. Thus, only the spectra in this frequency range were used in our study. For each cycle, the temperature corresponding to the measured power is determined by differencing the spectral line data acquired with open and closed noise diode, followed by calibrating the resulting power values of all spectral lines into temperature. The calibrated spectral data are processed using the astronomical data processing software GILDAS/CLASS222https://www.iram.fr/IRAMFR/GILDAS/. For all observations, a second-order polynomial baseline was fitted and subtracted over a velocity span of 60 km s-1 around the line, excluding the channels containing line emission, absorption features, or identifiable RFI.
@@
We performed Gaussian fitting on all spectra with clear detections to obtain the integrated intensity, line-center velocity, FWHM, and peak amplitude of the OH 18-cm lines. We also applied a trapezoidal fitting procedure to the 1667 MHz spectra of the four observations with detections October 23, 26 and November 4, 5 in order to derive the half-width of the trapezoidal base. The derived parameters are summarized in Table 3.
@@
| Date (UT) | Lines | RMS (K) | Area (K km s-1) | (K) | S ( Jy km s-1) | (km s-1) | FWHM (km s-1) | (K) | (km s-1) | ( s-1) | |
| 2025/10/23 | 1665 1667 | 0.07 0.07 | 2.10.1 3.90.1 | 0.45 | 3.3 | 1.70.1 3.10.1 | 13.40.1 13.30.1 | 3.40.2 3.10.1 | 0.59 1.21 | 2.40.3 2.40.1 | 1.10.1 |
| 2025/10/26 | 1665 1667 | 0.06 0.06 | 2.30.1 4.80.1 | 0.51 | 3.5 | 1.80.1 3.80.1 | 28.80.1 29.00.1 | 3.50.2 3.40.1 | 0.61 1.33 | 2.90.6 2.60.1 | 1.30.1 |
| 2025/11/03 | 1665 1667 | - 0.09 | - -0.40.1 | -0.27 | 3.1 | - -0.40.1 | - 53.50.1 | - 1.10.2 | - -0.34 | - | - |
| 2025/11/04 | 1665 1667 | - 0.09 | - -1.40.2 | -0.37 | 3.1 | - -1.30.2 | - 56.40.2 | - 3.20.4 | - -0.40 | - 3.60.8 | 1.50.3 |
| 2025/11/05 | 1665 1667 | - 0.09 | - -1.30.2 | -0.33 | 3.1 | - -1.40.2 | - 57.00.3 | - 3.90.6 | - -0.33 | - 3.90.9 | 1.70.4 |
Notes. Column 2 represents the spectral lines observed in the corresponding frequency range; Column 3 is the 1 noise of the observed spectra with the channel width 0.38 km s-1; Column 4, 8, 9, 10 are derived from Gaussian fitting, representing the integrated intensity, the position of the center of the line obtained through fitting, the full width at half height, and the peak temperature obtained through fitting, respectively; Column 5 is the inversion of ground state of (Schleicher and A’Hearn 1988); Column 7 represents the corresponding flux; Column 11 is the half-length of the base, derived from trapezoidal fitting which represents the sum of the velocities of the parent and daughter molecules; Column 12 is the derived water production rates of this comet.
@@ 4 Results and Discussion
During four observing epochs on October 23, 26 and November 4, 5, either emission or absorption of the 1667 MHz line was detected.
@@ 4.1 Change of OH 18-cm Lines from Emission to Absorption
In our observations, the OH 18-cm lines exhibit a clear evolution from emission to absorption. During the first two observing epochs, the OH lines are detected in emission in both the 1665 and 1667 MHz transitions, with corresponding line-center velocities of 13.4 km s-1 and 28.8 km s-1, respectively. In subsequent observations, the lines appear in absorption in both main lines, with a stronger signal at 1667 MHz. The corresponding line-center velocities are 53.5 km s-1, 56.4 km s-1, and 57.0 km s-1 on November 3 to 5, respectively. The fitted line-center velocities are consistent with the expected topocentric radial velocities derived from JPL horizons system at each observing epoch, confirming the cometary origin of the detected signals. No distinct OH signal is detected during the observations from November 6 to 8 within the expected velocity range.
@@
such that at Å, a heliocentric radial velocity of corresponds to a wavelength of Å. At these wavelengths, the solar Fraunhofer spectrum is strongly structured on scales of order 0.1 Å, comparable to the Doppler shifts induced by heliocentric radial velocities of a few km s-1 (see Fig. 2 of Schleicher and A’Hearn 1982).
Consequently, modest heliocentric radial velocity variations during the observing period can alter the OH pumping rate and drive the inversion parameter of the ground-state -doublet across zero. Against the Galactic background continuum, this leads to a transition in the observed OH 18-cm line profiles from emission to absorption (Despois et al. 1981). Using the inversion curves of Schleicher and A’Hearn (1988), we derived the inversion parameter at the heliocentric radial velocity of each epoch and compared it with the observed line profiles. The emission detected at early epochs and the absorption observed at later times are consistent with positive () and negative () inversion, respectively. The absence of a clear OH detection in the final observing epochs might be attributed to several factors. One possible reason is that the inversion parameter is close to zero, which results in a very weak OH signal. Another possibility is that the geocentric distance increased by a factor of 1.7, from 0.60 AU on October 23 to 1.01 AU on November 8, as the comet was moving away from the Earth. Consequently, the OH intensity may have fallen below the sensitivity limit of the telescope, leading to a non-detection.
@@ 4.2 Water Expansion Velocities from OH Lines
The OH coma is commonly described using a spherically symmetric, steady-state, and isotropic outflow model, in which OH is treated as a secondary product of H2O photodissociation. The velocity of OH molecules is determined by the expansion velocity of the parent gas, , combined with an additional ejection velocity due to photodissociation, (Bockelee-Morvan and Gerard 1984; Crovisier et al. 2002). The OH 18-cm lines profiles can be well reproduced by trapezoidal fitting, where the width of the lower base corresponds to 2(+). This trapezoidal profile model has been widely applied in cometary OH line studies (Drozdovskaya et al. 2023; Li et al. 2025; Sakai et al. 2025). Although OH line profiles always display asymmetries due to anisotropic outgassing toward the Sun or Earth (Li et al. 2025), such effects have only a limited impact on the determination of the lower-base width (Bockelee-Morvan and Gerard 1984). In this work, we adopt a value of = 0.95 km s-1 to represent the ejection velocity of OH radicals following the photodissociation of H2O, based on the analyses presented in (Crovisier et al. 2002; Tseng et al. 2007). This value is consistent with the typical OH ejection speeds reported in previous cometary studies (Crovisier 1989; Bockelée-Morvan et al. 1990). The resulting values of derived from the symmetric trapezoidal fits are listed in Table 3.
@@ 4.3 The estimate of OH production rate
where is the total number of OH radicals in the coma, is the integrated intensity of the 1667 MHz line in units of Jy km s-1, is the inversion of the ground state, and is the geocentric distance of the comet in AU. denotes the background temperature in the unit of K, is the OH lifetime, and is correction factor. For , it can be expressed as: = (Wang et al. 2017; Sakai et al. 2025), and was estimated as s from Bockelée-Morvan et al. (1990), where the is the heliocentric distance of the comet. Following the method of Crovisier et al. (2002), we estimated using the 1420 MHz brightness temperature distribution. The 1420 MHz map was obtained from the NASA/LAMBDA333https://lambda.gsfc.nasa.gov/product Stockert–Villa-Elisa all-sky survey (Reich 1982; Reich and Reich 1986; Testori et al. 2001), and the corresponding background temperature is listed in Table 3. The inversion of ground state at the corresponding coma velocity is obtained from (Schleicher and A’Hearn 1988) and listed in Table 3.
@@
where is the OH column density distribution derived from Haser model. The expansion velocity is derived from the trapezoidal fits, and the photodissociation lifetime of H2O at 1 AU is adopted as s (Huebner et al. 1992), also scaling as . The beam weighting function is assumed to be Gaussian, with a FWHM of 2.5 arcmin.
@@
When only the beam correction is applied, the derived values serve as lower limits, with value of , , , and s-1 for October 23, 26 and November 4, 5, respectively. The production rate basically increases as the heliocentric distance decreases. The water production rate of comet can be approximated as (e.g., Sakai et al. 2025; Li et al. 2025).